OpenStudy (anonymous):

A theater had 100 people in the audience consisting of men, women, and children. The ticket prices were: men - $10 each; women - $3 each; children - $0.50 each. The theater collected a total of $100 from 100 people. How many men, women, and children were in the audience?

OpenStudy (anonymous):

Systems of equations! You have 100 people total, so you have m+w+c=100 Where m,w,c are the number of men/women/children respectively. Also, 10m+3w+.5c=100. Try solving for m,w,c.

OpenStudy (anonymous):

m+w+c=100 10(m) + 3(w) + .5(c) = 100(100) we have 2 equations, but 3 unknown variables: m, w, c. there isn't anyway to obtain an answer that has a value, is there any more information?